# Less is More: Skim Transformer for Light Field Image Super-resolution

### v2

@@ Abstract
A light field image captures scenes through an array of micro-lenses, providing a rich representation that encompasses spatial and angular information. While this richness comes at the cost of significant data redundancy, most existing light field methods still tend to indiscriminately utilize all the information from sub-aperture images (SAIs) in an attempt to harness every visual cue regardless of their disparity significance. However, this paradigm inevitably leads to disparity entanglement, a fundamental cause of inefficiency in light field image processing. To address this limitation, we introduce the Skim Transformer, a novel architecture inspired by the “less is more” philosophy. Unlike conventional light field Transformers, our Skim Transformer features a multi-branch structure where each branch is dedicated to a specific disparity range by constructing its attention score matrix over a skimmed subset of SAIs, rather than all of them. Building upon this core component, we present SkimLFSR, an efficient yet powerful network for light field super-resolution (LFSR). Requiring only 67% of parameters, SkimLFSR achieves state-of-the-art results surpassing the best existing method by an average of 0.59 dB and 0.35 dB in PSNR at the 2×2$\times$2 × and 4×4$\times$4 × tasks, respectively. Through in-depth analyses, we reveal that SkimLFSR, guided by the predefined skimmed SAI sets as prior knowledge, demonstrates distinct disparity-aware behaviors in attending to visual cues. These findings highlight its effectiveness and adaptability as a promising paradigm for light field image processing.
Index Terms:
Light field, Super-resolution, Image processing, Deep learning.
@@ I Introduction
Light Field (LF) imaging captures light rays of a scene from multiple angular directions in a single shot, enriching the scene representation and enabling advanced computer vision capabilities such as post-capture refocusing [1, 2], depth estimation [3, 4, 5, 6, 7], material recognition [8, 9], salient object detection in complex scenes [10, 11, 12], microscopy [13, 14, 15], anti-spoofing face recognition [16, 17], and virtual reality (VR) media enhancement [18, 19]. The development of LF acquisition devices spans from early laboratory-built camera arrays [20, 21] to practical cameras, such as Raytrix [22] and Lytro Illum [23], and the more recent panoramic LF system [19] built by Google.
Due to technical limitations, accommodating a large number of micro-lenses within a compact form factor often comes at the cost of spatial resolution in LF images. For instance, the spatial resolution of a Lytro Illum camera is down to 376×541376$\times$ 541376 × 541. To mitigate this limitation, several light field image super-resolution (LFSR) methods have been proposed to enhance the spatial resolution while preserving LF’s intrinsic angular structure.
Due to the nature of LF imaging, the angular structure is inherently non-uniform and can vary significantly across different spatial locations and angular directions. These variations arise primarily from two factors: scene depth (from an image-wise perspective) and camera configuration (from a device perspective). Disparity modeling plays a crucial role in harnessing this angular information to achieve effective LFSR.
Fig. 1 exemplifies how parallax manifests differently across disparity ranges. Specifically, Fig. 1(b) demonstrates a larger disparity range by computing the parallax between two outer sub-aperture images (SAIs), making foreground objects such as the Lego studs (highlighted in the blue box) appear more visually prominent. In contrast, Fig. 1(c), which reflects a smaller disparity range using inner SAIs, renders these structures more blurred and indistinct. Conversely, background elements like the brick wall edges (highlighted in the red box) exhibit excessive parallax in Fig. 1(b), whereas they appear sharper and more coherent in Fig. 1(c). These disparities reveal nuanced variations and complexity of disparity distribution in LF images, meanwhile, offering valuable cues for enhancing LFSR performance.
@@
Figure 1: Parallax at different disparity ranges in the Lego Knights sample. The bottom-right indicator in (b) and (c) indicates the SAIs used to obtain the parallax.
@@
The recent integration of deep learning techniques, especially Vision Transformers (ViT) techniques [24, 25, 26, 27], has demonstrated promising potential in leveraging the self-attention mechanism to establish long-range dependencies for spatial, angular, and epipolar plane image (EPI) subspaces in LF images. However, existing Transformer-based methods commonly overlook the substantial disparity variations inherent in LF data. These approaches typically apply self-attention uniformly across the entire LF tensor in a single pass, attempting to exploit all available cues without adequately considering their significance, instead relying on the self-attention mechanism to learn the relevance solely from training data.
In this paper, we identify this issue as disparity entanglement, where heterogeneous disparity cues are processed homogeneously. This not only impairs a LFSR network’s ability to model disparity but also introduces computational redundancy.
@@
To address this issue, we propose the Skim Transformer, a novel Transformer architecture grounded in the “less is more” philosophy. Rather than indiscriminately processing all SAIs, the Skim Transformer selectively samples a subset of SAIs, referred to as a skimmed SAI set, that serves as prior knowledge to effectively manage disparity information and guide the construction of self-attention. Meanwhile, a multi-branch structure is constructed to broaden this capability to handle diverse disparity ranges. As a result, the Skim Transformer processes less information, but not only mitigates computation redundancy but also disentangles disparity to extract more relevant information.
@@
Building upon Skim Transformers, we present SkimLFSR, an efficient and effective LFSR network that not only delivers superior performance but also improves efficiency through its paradigm of disparity disentanglement. In our experiments, a lightweight variant of SkimLFSR outperforms nearly all existing methods [25, 24, 25] while using only 37% of the parameters, 35% of the FLOPs, and 28% of the inference time. The full version of SkimLFSR surpasses the previous state-of-the-art [27] by a substantial margin of 0.59 dB and 0.35 dB in PSNR for the 2×2$\times$2 × and 4×4$\times$4 × LFSR tasks, respectively, while requiring just 67% of the parameters used by the prior leading method.
@@
Furthermore, we conduct a series of in-depth analyses to uncover the inner workings of SkimLFSR. The results reveal that SkimLFSR’s superior performance originates from its disparity-aware characteristic, as evidenced through deep feature analyses where SkimLFSR exhibits latent discriminative capability toward scene depths and camera configurations. Remarkably, this capability emerges even though neither of these two modalities is explicitly provided as supervision during the regression-based LFSR training, highlighting the effectiveness of Skim Transformer to implicitly capture and exploit underlying disparity cues.
@@
    We propose the Skim Transformer, a novel architecture built on the “less is more” philosophy. By selectively sampling a skimmed SAI set in its self-attention mechanism and adopting a multi-branch structure, Skim Transformer explicitly targets different disparity ranges and achieves disparity disentanglement in its disparity modeling process.
@@
    Built upon this, we present SkimLFSR, an efficient and high-performing LFSR network. It surpasses the previous leading method by 0.59 dB and 0.35 dB in PSNR for 2×2$\times$2 × and 4×4$\times$4 × LFSR, respectively, while requiring just 67% of the parameters used by the prior leading method.
@@ II Related Work
II-A Single Image Super-resolution
Single image super-resolution (SISR) is a foundational task in low-level computer vision, aiming to reconstruct a high-resolution (HR) image from a given low-resolution (LR) input. As an upstream research area, advancements in SISR often contribute transferable insights and methodologies to other related domains, including LFSR.
The integration of deep learning into SISR was first introduced by Dong et al. [28] in their SRCNN, the first SISR network based on a convolutional neural network (CNN). It significantly outperformed traditional methods [29, 30], establishing a new paradigm for image processing.
Progress in SISR has since accelerated through a series of architectural innovations. VDSR [31] introduced very deep networks with residual learning to improve data flow in network propagation. RDN [32] enhanced hierarchical feature fusion through dense connections. RCAN [33] further improved representational capacity by incorporating a residual-in-residual structure along with channel attention mechanisms.
Beyond spatial enhancement, researchers have explored the incorporation of auxiliary modalities to enrich the reconstruction process, including spectral priors [34] and language-based guidance via text-to-image models [35]. A distinct line of research has focused on improving perceptual quality, moving beyond traditional pixel-wise fidelity metrics like Pixel-wise Signal-to-Noise Ratio (PSNR). Key developments in this area include the introduction of perceptual loss functions and adversarial learning frameworks [36, 37].
More recently, Transformer-based architectures have gained traction in SISR, motivated by the success of Vision Transformers (ViTs) [38] in high-level vision tasks. Chen et al. [39] proposed a unified IPT Transformer backbone pre-trained across multiple image restoration tasks, demonstrating strong generalization across low-level tasks. SwinIR [40], based on the Swin Transformer [41], introduced a hierarchical design with shifted windows to effectively model both local and global dependencies. HAT [42] extended this paradigm by integrating channel attention with window-based self-attention to balance fine-grained detail preservation with contextual awareness.
Despite their superior performance, Transformer-based SISR methods are inherently constrained by the quadratic complexity of global self-attention with respect to input resolution, limiting their scalability. To overcome this bottleneck, efficient variants such as SRFormer [43] and ELAN [44] have been proposed. SRFormer employs a permuted self-attention, while ELAN introduces a number of modifications to the self-attention mechanism, including a shifted convolution and sharing attention scores across modules.
The recent emergence of natural language processing (NLP) has also inspired the development of language-assisted SISR methods. For instance, Sun et al. [45, 46] bridged the gap between text and image domains by leveraging diffusion [47], image-to-text [48, 49], and text-to-image [50] models.
Different from SISR, which primarily focuses on restoring visual details degraded during downsampling, LFSR imposes an additional constraint: the preservation and enhancement of angular consistency across SAIs. Moreover, the computation and memory constraints of LFSR are more severe than those of SISR, due to the higher dimensionality of LF data. These unique challenges motivate the development of dedicated LFSR methods.
II-B Light Field Image Super-resolution
Processing 4D LF images introduces substantial challenges for neural network design due to their high dimensionality and the intricate spatial-angular dependencies they encode. A naive solution is to apply 4D convolutions directly. However, it incurs prohibitively intensive computation, making both training and inference impractical, and it often suffers from optimization difficulties due to the overwhelming data complexity [51].
To alleviate these issues, decomposition-based strategies have gained prominence in the LF community. Wang et al. [8] introduced the interleaved filter for LF material recognition, which approximates 4D convolutions by decomposing them into separate spatial and angular convolutions. This technique achieved comparable performance to full 4D convolutions while substantially reducing computational overhead, laying the groundwork for efficient LF representation learning.
@@ II-B Light Field Image Super-resolution
Given that LF images inherently encompass multiple subspaces, the decomposition paradigm is extended beyond the spatial and angular subspaces. Cheng et al. [58] proposed LFSSR-SAV, which employs spatial-angular versatile convolution to handle EPI subspaces. Hu et al. [59] generalized this idea with the Decomposition Kernel Network (DKNet), jointly modeling spatial, angular, and EPI subspaces. Wang et al. [60] introduced a disentanglement mechanism to better aggregate and refine multi-subspace features, while Duong et al. [61] proposed a hybrid model that integrates multi-orientation EPI extractors with angular and spatial representation extractors.
@@
Parallel to advancements in SISR, the ViT technique has expanded its successes to LFSR. Wang et al. [62] proposed DPT equipped with content and gradient Transformers to build long-range dependencies within the spatial subspace. Liang et al. [26] employed self-attention on overlapped local windows in their spatial Transformers, akin to HAT [42] for SISR. Transformers are also extended to the EPI subspaces in EPIT [24]. Cong et al. proposed LF-DET [25], integrating a sub-sampling spatial strategy and a multi-scale angular modeling strategy to enhance spatial and angular modeling. Hu et al. [27] proposed M2MTNet with Many-to-Many Transformers that avoid the practice of merging subspaces into the batch dimension and resolve the broadly existing issue of subspace isolation.
@@
Despite these recent advancements, as discussed in the previous section, a key limitation remains: most existing methods treat all the LF clues equally within attention modules, leading to disparity entanglement. To address this issue, we propose the Skim Transformer, a novel disparity-aware architecture that selectively attends to relevant information.
@@ III-A Preliminaries
The task of LFSR is to enhance the spatial resolution of a low-resolution (LR) LF image $I_{LR}$ to produce a super-resolved (SR) LF image $I_{SR}$, which aims to approximate the corresponding high-resolution (HR) LF image $I_{HR}$:
@@ III-B Network Architecture
The proposed SkimLFSR network is illustrated in Fig. 2. The network comprises three primary stages: initial feature extraction, deep feature extraction, and image generation. The initial feature extraction stage utilizes four consecutive $\times 3$ 3 convolution layers operating on the spatial subspace of LF images to obtain low-level features. The subsequent deep feature extraction stage incorporates a sequence of $N_{CB}$ correlation blocks to derive comprehensive correlation information and establish a high-level spatial-angular representation. Their details will be elaborated on in the following paragraphs. The final image generation stage aggregates the deep features via convolution layers and upsamples the spatial resolution through a pixel shuffler [63] to produce the final LF image output.
@@
These Transformers are preceded by two $\times 3$ 3 spatial convolutions serving as positional encoding [64], and followed by a channel attention mechanism [33] for channel-wise enhancement.
Connection Enhancement. To improve SkimLFSR’s data flow, besides a standard skip connection that bridges the SR and LR, two extra connections are integrated, as illustrated in Fig. 2. Firstly, in alignment with the strategy from DKNet [59], a raw image connection concatenates the raw image tensor directly to the output of the last correlation block prior to the image generation stage. This connection acts as a special form of dense connections [65], ensuring that raw image data directly contributes to feature aggregation and upsampling. Secondly, inspired by RGT [66], each correlation block incorporates a learnable skip connection that dynamically adjusts the input by channel-wise coefficients $\alpha\in\mathbb{R}^{C}$ and adds it to the output. These two connections incur negligible computational overhead but usually improve the SkimLFSR’s performance by about 0.1 dB PSNR in the experiments.

### v3

@@ Abstract
A light field image captures scenes through its micro-lens array, providing a rich representation that encompasses spatial and angular information. While this richness comes at significant data redundancy, most existing methods tend to indiscriminately utilize all the information from sub-aperture images (SAIs) in an attempt to harness every visual cue regardless of their disparity significance. However, this paradigm inevitably leads to disparity entanglement, a fundamental cause of inefficiency in light field image processing. To address this limitation, we introduce the Skim Transformer, a novel architecture inspired by the “less is more” philosophy. It features a multi-branch structure where each branch is dedicated to a specific disparity range by constructing its attention score matrix over a skimmed subset of SAIs, rather than all of them. Building upon it, we present SkimLFSR, an efficient yet powerful network for light field image super-resolution. Requiring only 67% of the prior leading method’s parameters, SkimLFSR achieves state-of-the-art results surpassing the best existing method by 0.63 dB and 0.35 dB PSNR at the $2\times$ and $4\times$ tasks, respectively. Through in-depth analyses, we reveal that SkimLFSR, guided by the predefined skimmed SAI sets as prior knowledge, demonstrates distinct disparity-aware behaviors in attending to visual cues. Last but not least, we conduct an experiment to validate SkimLFSR’s generalizability across different angular resolutions, where it achieves competitive performance on a larger angular resolution without any retraining or major network modifications. These findings highlight its effectiveness and adaptability as a promising paradigm for light field image processing.
@@ I Introduction
Light Field (LF) imaging captures light rays of a scene from multiple angular directions in a single shot, enriching the scene representation and enabling advanced computer vision capabilities such as post-capture refocusing [13], depth estimation [18, 43], material recognition [44, 34], salient object detection [4], microscopy [41, 26], and virtual reality (VR) media enhancement [2]. The development of LF acquisition devices spans from early laboratory-built camera arrays [11, 39] to practical cameras, such as Raytrix [36] and Lytro Illum [50], and the more recent panoramic LF system [2] built by Google.
Due to technical limitations, accommodating a large number of micro-lenses within a compact form often comes at the cost of spatial resolution in LF images. For instance, the spatial resolution of a Lytro Illum camera is down to $376\times 541$. To mitigate this limitation, several light field image super-resolution (LFSR) methods have been proposed to enhance the spatial resolution while preserving LF’s intrinsic angular structure. Inherently, the LF’s angular structure is non-uniform and can vary significantly across different spatial locations and angular directions. These variations arise primarily from two factors: scene depth (from an image-wise perspective) and camera configuration (from a device perspective). Disparity modeling plays a crucial role in harnessing this angular information to achieve effective LFSR.
Fig. 1 exemplifies how parallax manifests differently across disparity ranges. Specifically, Fig. 1 (b) demonstrates a larger disparity range by computing the parallax between two outer sub-aperture images (SAIs), making foreground objects such as the Lego studs (the blue box) appear more visually prominent. In contrast, Fig. 1 (c), which reflects a smaller disparity range using inner SAIs, renders these structures more blurred and indistinct. Conversely, background elements like the brick wall edges (the red box) exhibit excessive parallax in Fig. 1 (b), whereas they appear sharper and more coherent in Fig. 1 (c). These disparities reveal nuanced variations and complexity of disparity distribution in LF images, meanwhile, offering valuable cues for enhancing LFSR performance.
@@
Figure 1: Parallax at different disparity ranges in Lego Knights. The bottom-right indicator in (b) and (c) indicates the SAIs for calculating the parallax.
@@
The recent integration of deep learning techniques, especially Vision Transformers (ViT) techniques [28, 9, 29, 21], has demonstrated promising potential in leveraging the self-attention mechanism to establish long-range dependencies for spatial, angular, and epipolar plane image (EPI) subspaces in LF images. However, existing Transformer-based methods commonly overlook the substantial disparity variations inherent in LF data. These methods typically apply self-attention uniformly across the entire LF tensor in a single pass, attempting to exploit all available cues without adequately considering their significance, instead relying on the self-attention mechanism to learn the relevance solely from training data.
@@
In this paper, we identify this issue as disparity entanglement, where heterogeneous disparity cues are processed homogeneously. This not only impairs a LFSR network’s ability to model disparity but also introduces computational redundancy. To address this issue, we propose the Skim Transformer, a novel Transformer architecture grounded in the “less is more” philosophy. Rather than indiscriminately processing all SAIs, the Skim Transformer selectively samples a subset of SAIs, referred to as a skimmed SAI set, that serves as prior knowledge to effectively manage disparity information and guide the construction of self-attention. Meanwhile, a multi-branch structure is constructed to broaden this capability to handle diverse disparity ranges. As a result, the Skim Transformer processes less information, but not only mitigates computation redundancy but also disentangles disparity to extract more relevant information.
@@
Building upon Skim Transformers, we present SkimLFSR, an efficient and effective LFSR network that not only delivers superior performance but also improves efficiency through its paradigm of disparity disentanglement. In our experiments, a lightweight variant of SkimLFSR outperforms nearly all existing methods [28, 9] while using only 37% of the parameters, 35% of the FLOPs, and 28% of the inference time. The full version of SkimLFSR surpasses the previous state-of-the-art [21] by a substantial margin of 0.63 dB and 0.35 dB in PSNR for the $2\times$ and $4\times$ LFSR tasks, respectively, while requiring just 67% of the parameters used by the prior leading method.
@@
We further conduct a series of in-depth analyses to uncover the inner workings of SkimLFSR. The results reveal that SkimLFSR’s superior performance originates from its disparity-aware characteristic, as evidenced through deep feature analyses where SkimLFSR exhibits latent discriminative capability toward scene depths and camera configurations. Remarkably, this capability emerges even though neither of these two modalities is explicitly provided as supervision during the regression-based LFSR training, highlighting the effectiveness of Skim Transformer to implicitly capture and exploit underlying disparity cues.
Finally, as Skim Transformers projects only a skimmed SAI set instead of the entire angular space into embedding, it is decoupled from the angular resolution that it is trained on, and as a result, it is inherently angular-resolution-agnostic, i.e., it can generalize to different angular resolutions. In the experiments, we will validate that Skim Transformers learn essential disparity information that is generalizable across different angular resolutions, which enables SkimLFSR to achieve competitive performance on an angular resolution that is different from the one it is trained on without any retraining and major network modifications.
@@
    We propose the Skim Transformer, a novel architecture built on the “less is more” philosophy. By selectively sampling a skimmed SAI set in self-attention and adopting a multi-branch structure, Skim Transformer explicitly targets different disparity ranges and achieves disparity disentanglement.
@@
    Built upon this, we present SkimLFSR, an efficient and high-performing LFSR network. It surpasses the previous leading method, while requiring substantially fewer parameters and shorter inference time, striking a good balance between performance and efficiency.
@@
  5. 5.
    We further discover that the Skim Transformer is inherently angular-resolution-agnostic, allowing SkimLFSR to seamlessly generalize to different angular resolutions and achieve very competitive performance without any retraining and major network modifications.
@@ II Related Work
Processing 4D LF images introduces substantial challenges for neural network design due to their high dimensionality and the intricate spatial-angular dependencies they encode. To alleviate these issues, Wang et al. [44] introduced the interleaved filter for LF material recognition, which approximates 4D convolutions by decomposing them into separate spatial and angular convolutions.
@@
Given that LF images inherently encompass multiple subspaces, the decomposition paradigm is extended beyond the spatial and angular subspaces. Cheng et al. [8] proposed LFSSR-SAV, which employs spatial-angular versatile convolution to handle EPI subspaces. Hu et al. [22] generalized this idea with the Decomposition Kernel Network (DKNet), jointly modeling spatial, angular, and EPI subspaces. Wang et al. [46] introduced a disentanglement mechanism to better aggregate and refine multi-subspace features, while Duong et al. [40] proposed a hybrid model that integrates multi-orientation EPI extractors with angular and spatial representation extractors. Zhang et al. [59] proposed a cascade residual learning based adaptive feature aggregation framework to aggregate features from the subspaces. This paradigm is also adopted in other LF-related tasks. Yeung et al. [55] proposed a spatial-angular separable convolution, consisting of a spatial convolution and an angular convolution, for LF angular super-resolution. In contrast to this task, typically performed on a single LF image with a limited field of view (FOV), Mao et al. [35] introduced a novel LF task called sparse-to-dense inbetweening, which aims to reconstruct a high–angular-resolution LF image from two low–angular-resolution inputs to extend the FOV. In their proposed method, the feature extraction process is decomposed into two EPI blocks and one spatial block.
Parallel to advancements in SISR, several techniques have been extended to LFSR, such as degradation estimation [1, 60], which has been incorporated into LFSR pipelines by Wang et al. [45] and Xiao et al. [54] to improve robustness against diverse degradation types. CutBlur [57] is a data augmentation strategy that copies patches from LR images and pastes them into HR images, encouraging the model to learn which regions require super-resolution rather than uniformly enhancing all pixels. Inspired by CutBlur, Xiao et al. [53] introduced a “cutting–blending–pasting” strategy that blends LR patches across different SAIs to enable the same capability in LFSR methods.
@@
The ViT architecture [12, 27, 5, 61] has been a popular architecture for SISR and has expanded its successes to LFSR. Wang et al. [42] proposed DPT equipped with content and gradient Transformers to build long-range dependencies within the spatial subspace. Liang et al. [29] employed self-attention on overlapped local windows in their spatial Transformers, akin to SwinIR [27] and HAT [5] for SISR. Transformers are also extended to the EPI subspaces in EPIT [28]. Cong et al. proposed LF-DET [9], integrating a sub-sampling spatial strategy and a multi-scale angular modeling strategy to enhance spatial and angular modeling. Hu et al. [21] proposed M2MTNet with Many-to-Many Transformers that avoid the practice of merging subspaces into the batch dimension and resolve the broadly existing issue of subspace isolation. Xiao et al. [52] incorporated SAI-wise occlusion maps as an additional branch and fused them with the initial visual features to achieve occlusion-aware LFSR.
More recently, Mamba-based methods [15, 10] have emerged as a promising alternative to Transformer-based methods. Its Selective State Space Models (SSM) uses input-dependent selective scanning to propagate a latent state across the sequence, and its computational complexity is linear with the sequence length, compared to the quadratic complexity of Transformer-based methods. This property has attracted the computer vision community to adopt it in various tasks [62, 16]. For LFSR, Xia et al. [51] proposed LFMamba, which adopts SSM to model the spatial, angular, and EPI subspaces as sequences. While the EPI subspace is for modeling spatial-angular correlations, Wei et al. [49] argue that EPI-based methods are limited by locality constraints because their information comes from a single epipolar line. They proposed L2FMamba, which adopts SSM to directly perform on individual SAIs, across SAIs, and on the macropixel image (MacPI). Liu et al. [30] proposed L3FMamba, which adopts a SSM-based method with dark, bright, and average channel prior injection for low-light LF image enhancement.
@@
Despite these recent advancements, as discussed in the previous section, a key limitation remains: most existing methods treat all the disparity cues indiscriminately, leading to disparity entanglement. Our Skim Transformer will be introduced to address this issue.
Figure 2: Illustration of the SkimLFSR network.
@@ III-A Preliminaries
The LFSR task aims to enhance the spatial resolution of a low-resolution (LR) LF image $I_{LR}$ and produce a super-resolved (SR) LF image $I_{SR}$ to approximate the corresponding high-resolution (HR) LF image $I_{HR}$:
@@ III-B Network Architecture
The proposed SkimLFSR network is illustrated in Fig. 2. The network comprises three primary stages: initial feature extraction, deep feature extraction, and image generation. The initial feature extraction stage utilizes four consecutive $3\times 3$ convolution layers operating on the spatial subspace of LF images to obtain low-level features. The subsequent deep feature extraction stage incorporates a sequence of $N_{CB}$ correlation blocks to derive comprehensive correlation information and establish a high-level spatial-angular representation. The final image generation stage aggregates the deep features via convolution layers and upsamples the spatial resolution through a pixel shuffler [38] to produce the final LF image.
@@
Connection Enhancement. To improve SkimLFSR’s data flow, besides a standard skip connection that bridges the SR and LR, two extra connections are integrated, as illustrated in Fig. 2. Firstly, in alignment with DKNet [22], a raw image connection concatenates the raw image tensor directly to the output of the last correlation block prior to the image generation stage. This connection acts as a special form of dense connections [17], ensuring that raw image data directly contributes to feature aggregation and upsampling. Secondly, inspired by RGT [7], each correlation block incorporates a learnable skip connection that dynamically adjusts the input by channel-wise coefficients $\alpha\in\mathbb{R}^{C}$. These two connections incur negligible computational overhead but empirically improve the network performance by about 0.1 dB PSNR.